# Imaging coupled vibrational, rotational, and electronic wave packet dynamics in a triatomic molecule

### v1

@@ Abstract
Molecular dynamics triggered by interaction with light often involve the excitation of several electronic, vibrational, and rotational states. Characterizing the resulting coupled electronic and nuclear wave packet motion represents a severe challenge, even for small polyatomic systems. In this Letter, we demonstrate how the interplay between vibrational, rotational, and electronic degrees of freedom governs the evolution of molecular wave packets in the low-lying states of strong-field-ionized sulfur dioxide. Using time-resolved Coulomb explosion imaging (CEI) in combination with quantum mechanical wave packet simulations, we directly map bending vibrations of the molecule, show how the vibrational wave packet is influenced by molecular alignment, and elucidate the role of the coupling between the two lowest electronic states of the cation. A conical intersection between these states couples the bending and asymmetric stretching coordinates, which is clearly reflected in the correlated fragment momenta. Our results suggest that multi-coincident CEI represents an efficient experimental tool for characterizing coupled electronic and nuclear motion in polyatomic molecules.
Recording molecular movies of an isolated molecule undergoing ultrafast structure changes with sufficient spatiotemporal resolution has been a longstanding goal in molecular optical sciences [1, 2, 3, 4]. Remarkable progress towards this goal has been demonstrated recently, including the development of gas-phase ultrafast electron diffraction (UED) and coherent x-ray imaging (CXI) techniques [5, 6, 7, 8]. While providing excellent insights into many photochemical reactions, UED still has limited time resolution, and CXI requires very bright x-ray sources due to its low-scattering cross-section. Both methods often need access to large-scale facilities; and the structural information is typically projected onto a single coordinate (i.e., the momentum transfer) which can cause difficulty in unraveling the complex dynamics of molecules. Directly observing atomic motions during chemical processes is still one of the grand challenges [2, 4].
Another method used for making such movies is time-resolved Coulomb explosion imaging (CEI) [9, 10, 11, 12, 13, 14, 15, 16]. Although CEI does not have the same strength in precise structure determination as UED or CXI, it provides excellent temporal resolution, high sensitivity to light atoms, and, most importantly, direct access to multidimensional information when coupled with coincidence imaging technique [17, 18, 19, 20]. Thus, it possesses an advantage in decoding and separating different reaction coordinates from the high information content caused by many entangled degrees of freedom. CEI is also a quasi-background-free technique allowing extraction of rare stochastic events (such as roaming) out of the measurement [21].
Light-induced vibrational wave packets play a fundamental role in many types of molecular dynamics and have been studied extensively (see, for instance, [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]). However, making such aforementioned movies for these processes is challenging due to its fast timescale and multimode complexity. Most experimental movies of molecular vibrations have been limited to diatomic molecules (such as [33, 19], [34, 35], , , [36], and [37, 38]) or the stretching of only one bond in polyatomic molecules such as [39]. Only a limited number of direct structural measurements of other vibrations in polyatomic molecules has been reported [40, 41, 42].
In this Letter, we demonstrate that laser-induced Coulomb explosion can directly image the ultrafast coherent wave packet of a laser-ionized small molecule, predominantly exhibiting bending vibrations influenced by rotational motion and electronic transitions. This technique produces a molecular movie showing the motion of individual atoms with sufficient temporal resolution. Our observations are supported by quantum mechanical wave packet calculations and classical Coulomb explosion simulations. Specifically, we investigate , a triatomic molecule which has been extensively studied over the last century (see, for instance, [43, 44, 18, 11, 45, 46, 47, 48, 49, 50, 51]) due to its significance in atmospheric processes and intriguing photodynamics, including the efficient formation of neutral or ionic molecular oxygen via various mechanisms and electronic states [52, 53, 54].
@@
This experiment utilized a Ti:Sapphire laser operating at 10 kHz, generating near-infrared (NIR) pulses with a central wavelength of approximately 790 nm and a pulse duration of about 28 fs. We split this output into pump and probe pulses, with independently controlled power, and scan their relative delay using a motorized linear stage. As sketched in Fig. 1, we use a strong laser pulse ( W/cm2) to ionize SO2 and launch vibrational wave packets in its cationic states. Subsequently, a second pulse of higher intensity ( W/cm2) is employed to further ionize and dissociate (Coulomb-explode) the molecules. The ionic products are detected in coincidence using a COLTRIMS apparatus [55]. The asymptotic momenta of these ions, determined from their measured time of flight and impact position, are used to deduce information about the induced wave packets.
@@
Previous photoionization studies [56, 57, 58, 59] have reported exclusive signatures of the bending progression of (00)+ in the SO () SO2 () photoelectron spectra. This is because the equilibrium geometries in the electronic ground state of the neutral SO2 () and ionic SO () have almost identical SO bond length but different OSO bond angle [56, 60], leading to a strong excitation into the bending vibrational mode () and very weak excitation into the symmetric ( and asymmetric () modes. The first six bending modes have very similar frequencies (within cm-1), approximately 400 cm-1. We found that the bending vibration influences many dynamics observed in our multi-dimensional coincidence data. For example, as shown in Fig. 1, the fast Fourier transform (FFT) power spectra from the delay-dependent yields of SO+ and SO show a strong signature of this bending vibration ( 400 cm-1 and the weak overtone at 800 cm-1). They also show a weaker but clear signature of the bending vibration in the ground state of neutral SO2 ( 520 cm-1) [61]. In this Letter, we focus on the three-body channel to directly visualize the bending vibration in the cation via delay-dependent kinetic energy (KE) spectra and angle correlations between the fragment ions.
@@
The main results from the 3-body coincidence channel are shown in Fig. 2. We present experimental data illustrating the delay dependence of several key observables: (a) the angle between momentum vectors of two O+ ions [denoted as ], (b) the kinetic energy of S+ fragment [denoted as KE()], and (c) the total kinetic energy of two O+ fragments [denoted as KEsum()]. Corresponding simulations of these observables are depicted in panels (d-f), and their mean versus delays are shown in panel (g). Panel (h) features a Newton plot at a specific delay of 250 fs, coinciding with the inner turning point of the bending wave packet.
@@
There are two distinguished features in the experimental data shown in Fig. 2(a-c). First, all the observables, namely , KE(), and KEsum(), show pronounced oscillations. The periodicity of these observed oscillations is 83 fs (corresponding to a frequency of 400 cm-1) clearly indicates that we mainly image the vibrational wave packet in the ground state of SO [57]. Second, there is a clear correlation between these observables: KE() oscillates out of phase while KEsum() oscillates in phase with [see Fig. 2(g)]. This can be explained using a simple picture: Due to the symmetry of the molecule, only vibrates along the symmetry axis; larger corresponds to lower KE() because the Coulomb forces from the two fragments on cancel each other out along the OO-axis more at larger angle, and vice versa.
Simulations corresponding to the experimental observables in Fig. 2(a-c) are presented in Fig. 2(d-f). We model the strong-field pump by solving the coupled-channel Schrödinger equation for the nuclear motion on Born-Oppenheimer potential energy surfaces of two electronic ground states of SO2 () and SO (), including all vibronic degrees of freedom (symmetric stretching, bending, and antisymmetric stretching modes) [63]. These two electronic states were calculated ab initio by using the multiconfigurational self-consistent-field method with seven frozen K-shell electrons and remaining electrons in 12 active orbitals, cc-pVTZ basis set within the GAMESS quantum chemistry code [64]. The wave packet induced by the pump is projected vertically to the triply-charged Coulombic potential (which is an approximation). We propagate this wave packet to large distances to simulate the explosion of the triply-charged parent molecule into the (O+, O+, S+) asymptote. We then Fourier transform the wave functions in real space to momentum space and calculate the experimentally measured observables as shown in Fig. 2(d-f). The simulation closely resembles the experimental data, manifesting nearly identical periodicity and phase behavior. The simulation can also explain the correlations between the momentum-space angle and kinetic energies of the fragment ions as a function of the real-space angle. Therefore, even though the quantitative relation between the measured momentum-space angle and the real-space bond angle is nontrivial and requires further analysis, these results already provided a direct and intuitive picture of the strong-field-induced vibrational wave packet in a triatomic molecule. While the observables for the triply-charged final state presented in Fig. 2 are dominated by the ionic ground-state vibrations, the FFTs of the delay-dependent data also shows a weak but clear signature of bending vibrations in the ground state of neutral SO2, similar to lower charge states (see Fig. 1 and SM, Fig. S2).
We further employ the strength of our multidimensional coincidence imaging measurement to investigate individual atoms and visualize their correlated atomic motions. We represent all ions in a momentum image (Newton plot) where each ion corresponds to an atom in the molecule. Fig. 2(h) features a representative image at 250 fs (the inner turning point of the bending wave packet). A timed sequence of these images visualizes the molecular structural changes as a function of delay, projected to the ion-momentum space by the Coulomb explosion probe. Such movies are provided in the SM, showing the correlated motion of individual atoms within a molecule as a function of time, directly reflecting the coherent bending vibration. In this specific case, the corresponding momentum image of each atom follows the same direction as its motion in real space (see SM, Fig. S3), giving a striking and intuitive picture of the nuclear motion. In other words, the manifests the same behavior as the OSO bond angle of the nuclear wave packet propagating in the ionic ground state. The movie also looks intuitive partly because, during the first 400 fs, the wave packet is well localized and mimics the motion of a “ball-and-stick” model.
@@
The total yield of the (O+, O+, S+) channel remains relatively flat and does not exhibit oscillations at the frequency of the bending vibration. This suggests that triple ionization is relatively insensitive to this mode. However, as depicted in Fig. 3(a), the yield initially decreases by approximately 15% within the first few hundred femtoseconds before flattening out. To elucidate this behavior, we examine the molecular rotation by plotting of the (O+, O+, S+) channel, depicted as the red curve in the same graph. Here, is the angle between the laser polarization and , representing the rotation of the OO-axis. The extracted rotation time scale is consistent with the two-body channels shown in SM, Fig. S4. The strong correlation between these two quantities (except near the overlap region) suggests that the observed yield variation can be attributed to the initial alignment of the pump-induced rotational wave packet. Although the OO-axis is the most polarizable axis, strong-field ionization from SO2 to SO is more likely to occur along the symmetry axis (see SM, Fig. S5, which includes Refs. [65, 66, 67, 68, 69]). Near t0, the symmetry axis tends to align with the laser polarization; subsequently, the molecules rotate, reaching initial alignment around 200 fs. At this time, the OO-axis is more likely to align with the laser polarization direction before returning to a more quasi-isotropic distribution.
This, in turn, partly explains the modulation of the mean shown as the green curve in Fig. 3(b). The pump-induced wave packet is a mixture of the neutral and ionic bending wave packets, where the ground ionic state has a larger equilibrium bond angle and thus contributes more events with larger . At the peak of the initial alignment, when the OO-axis is aligned along the laser polarization, the probe is likely to ionize more molecules from the ionic state (see SM, Fig. S6), resulting in the mean peaking near 200 fs.
Our results on direct imaging of the bending vibration also provide valuable insights into other phenomena, such as O formation in the SO, as depicted in Fig. 3(b). The strong anticorrelation between the (green) and yield of the (S+,O) channel (purple) offers an intuitive mechanistic explanation of the O yield modulation from the structure perspective. When the two oxygen atoms are further apart at the outer turning point [large ], the probability of producing O with the probe pulse decreases. Conversely, when the two oxygen atoms are closer at the inner turning point [small ], the likelihood of O formation with the probe pulse increases.
Finally, by comparing the KE sharing between the two O+ ions [see Fig. 4(a,b)], we observe that events with large (ranging from 140∘ to 150∘) exhibit symmetric KE sharing. In contrast, a significant number of events with small (between 90∘ and 100∘) displays asymmetric KE sharing. We attribute this feature to cations transitioning from the ground SO () to the first excited ionic state SO () by resonant one-photon transition. These two states have a symmetry-allowed conical intersection (CI) with the branching space comprised of the bending and asymmetric stretching coordinates [48]. After excitation to the SO () state near the inner turning point of the ground state bending mode, the wave packet travels towards the CI at , where it is coupled to the asymmetric stretch mode. This resonant one-photon dipole coupling can be induced either by the trailing edge of the pump pulse or the leading edge of the probe pulse before multiple ionization. The population in the state is predicted to quickly return to the ground state within less than 20 fs with only weak revival near 100 fs [48]. Fig. 4(c) shows that events with symmetric (blue) and asymmetric (green) KE sharing [all with small ] have distinct delay-dependent yields. Events with asymmetric KE sharing sharply peak at t0 and drop within less than 30 fs. Events with symmetric KE sharing (which could get to small with or without the coupling) do not have a sharp enhancement at t0 and are strongly modulated by the bending motion of the SO () state. The in-phase oscillation of the two curves in Fig. 4(c) is consistent with population transfer to the SO () state near the inner turning point of the ground state bending vibration. To further test this hypothesis of resonant inter-state coupling, we use a single laser pulse at 405 nm to ionize SO2 into (O+, O+, S+) and compare the KE sharing of the two O+ ions in this case with the 790 nm case. This comparison is shown in Fig. 4(d). The absence of asymmetric KE sharing (side peaks) in the 405 nm case suggests a resonant coupling near 790 nm, corroborating our conclusion that we see signatures of mode-coupling at the CI.
In conclusion, our work highlights the direct sensitivity of CEI to the change in the spatial distribution of atoms, demonstrating its feasibility in tracking structural changes by monitoring the correlated motion of individual atoms within a molecule. This approach offers intuitive mechanistic insights into various molecular dynamics. The high-dimensional measurements provide rich information, allowing us to observe and interpret the interplay between different degrees of freedom—such as rotation, vibration, and electronic motion—how they influence the evolution of the molecular wave packet, and determine experimental outcomes. Recent advancements have shown that CEI can image detailed 3D structures of gas-phase intermediate-sized molecules with about ten atoms [70, 71]. This capability opens the potential to follow structural changes in more complex dynamics of chemically relevant systems at similar levels of detail [72, 73, 74]. Our simulation captures the periodicity and phase behavior of different observables and their correlations, strongly supporting our interpretations. However, the quantitative connection between momentum-space and real-space observables remains a challenge. We plan to address this via improved modeling of the probe step.

### v3

# Simultaneous imaging of vibrational, rotational, and electronic wave packet dynamics in a triatomic molecule

@@ Abstract
Light-induced molecular dynamics often involve the excitation of several electronic, vibrational, and rotational states. Since the ensuing electronic and nuclear motion determines the pathways and outcomes of photoinduced reactions, our ability to monitor and understand these dynamics is crucial for molecular physics, physical chemistry, and photobiology. However, characterizing this complex motion represents a significant challenge when different degrees of freedom are strongly coupled. In this Letter, we demonstrate how the interplay between vibrational, rotational, and electronic degrees of freedom governs the evolution of molecular wave packets in the low-lying states of strong-field-ionized sulfur dioxide. Using time-resolved Coulomb explosion imaging (CEI) and quantum mechanical wave packet simulations, we directly map the bending vibrations of the molecule, show how the vibrational wave packet is influenced by molecular alignment, and elucidate the consequences of nuclear motion for the coupling between the two lowest electronic states of the cation. Our results demonstrate that multi-coincident CEI can be an efficient experimental tool for characterizing coupled electronic and nuclear motion in polyatomic molecules.
Recording “molecular movies” — tracking the motion of individual atoms during ultrafast structural changes with sufficient spatiotemporal resolution — has been a long-standing goal in molecular optical sciences [1, 2, 3, 4]. Remarkable progress toward this goal has been demonstrated recently, largely driven by the development of gas-phase ultrafast electron diffraction (UED) [5, 6] and ultrafast X-ray scattering (UXS) [7, 8, 9], which are particularly sensitive to nuclear motion. Together with spectroscopic methods sensitive to the electronic structure and advanced theoretical modeling, these techniques have provided critical insights into several important photochemical reactions [10, 2, 6]. However, despite the variety of available experimental tools, complete and unambiguous characterization of molecular dynamics remains challenging. Central to these challenges is the ubiquitous interplay between nuclear and electronic motion, which underpins essential processes in photochemistry, photophysics, and photobiology, including energy transfer, photodissociation, and nonradiative relaxation [11]. Such coupled motions are inherently complex, even for small polyatomic molecules [12]. Electronic excitations reshape molecular geometries and populate a variety of vibronic levels, while nuclear-electronic couplings enable nonadiabatic transitions near conical intersections, where electronic states converge [13].
Time-resolved Coulomb explosion imaging (CEI) is another method used to create “molecular movies” [14, 15, 16, 17, 18, 19, 20]. While CEI does not directly yield real-space molecular geometries, it provides excellent temporal resolution, high sensitivity to light atoms, and accessibility in university-scale laboratories. Critically, unlike UED and UXS, which project structural information onto a single coordinate (momentum transfer), coincidence-mode CEI [21, 22, 23, 24, 25, 26, 27] generates high-dimensional data with naturally embedded correlations that enable reliable separation of reaction pathways, even for rare stochastic processes [28]. Previously, both CEI and diffraction-based techniques have been employed to directly image light-induced rotational [29, 15, 30, 31, 19] and vibrational [23, 32, 33, 34, 8] wave packets, which play a critical role in many types of molecular dynamics [35, 36, 37, 38, 39, 40, 41, 42, 43]. However, most experimental movies of molecular vibrations have been limited to diatomic molecules (such as [44, 23], [45, 46], [47], [48, 47], [49, 47], and [32, 33, 34]) or the stretching of just one bond in polyatomic molecules [50, 51, 52]. Due to the fast time scale and complexity of rovibronic motions in polyatomic molecules, only a limited number of direct structural measurements of other vibrational modes has been reported [53, 54, 8].
In this Letter, we demonstrate that laser-induced Coulomb explosion can directly image ultrafast coherent wave packets in a small polyatomic molecule and produce movies that show the motion of individual atoms even in the presence of coupled nuclear and electronic motion. We investigate sulfur dioxide (SO2), a bent triatomic molecule that has been studied extensively (see, for instance, [55, 56, 22, 16, 57, 58, 59, 60, 61, 62, 63]) due to its significance in atmospheric processes and its intriguing photochemistry, including the efficient formation of neutral and ionic molecular oxygen [64, 65, 66, 67]. We employ time-resolved coincident CEI to create a movie of the coherent vibrational wave packet in SO and show that its dynamics are significantly influenced by the rotational motion. Moreover, we find that a secondary wave packet resulting from a resonant dipole coupling between the two lowest electronic states of SO traverses a conical intersection and leaves clear signatures of the nonadiabatic coupling between these two states in the ion momentum distributions.
@@
Our experiment utilized a Ti:Sapphire laser operating at 10 kHz, generating near-infrared (NIR) pulses with a central wavelength of approximately 790 nm and a pulse duration of about 28 fs. We split this output into pump and probe pulses, with independently controlled power, and scan their relative delay using a motorized linear stage. As depicted in Fig. 1, we use a 3.4 W/cm2 pump pulse to ionize SO2 and a 8 W/cm2 probe pulse to further ionize and dissociate (i.e., Coulomb explode) it. The ions are then detected in coincidence using a COLTRIMS apparatus [68]. The momenta of these ions, determined from their measured time of flight and impact position, are used to deduce information about the induced wave packets.
@@
Previous studies [69, 70, 71, 72] have reported that the bending mode progression dominates the SO () SO2 () photoelectron spectra because the equilibrium geometries in these electronic ground states have almost identical bond lengths but different bond angles [69, 73]. This leads to a strong excitation into the bending mode () and very weak excitation into the symmetric ( and asymmetric () stretch modes. The first seven bending-mode energies are approximately equally spaced with a nearest-neighbor separation of cm-1. In our experiment, which has a 17 cm-1 frequency resolution, we see oscillations at this frequency (and its first overtone at cm-1) in different observables for various ionization and fragmentation channels. This indicates that these channels involve the population of the SO ground state by the pump pulse and subsequent ionization and/or fragmentation by the probe. For example, the fast Fourier transform (FFT) spectra of the delay-dependent yields of SO+, SO and channels show a strong peak at the bending frequency and a much smaller one at the overtone (Fig. 1). They also show a weak but clear signature of the bending vibration in the ground state of neutral SO2 (at 520 cm-1) [74]. The latter could be induced by Raman excitation (analogous to bond softening), or by geometry-dependent ionization (so-called “Lochfrass” process) [75, 45, 76, 51, 77]. The determination of the relative importance of these mechanisms requires a detailed analysis of the wave packet phase and will be discussed elsewhere. Here, we focus on the three-body channel to directly visualize the bending vibrations of the cation via delay-dependent kinetic energy (KE) spectra and angle correlations between the fragment ions.
@@
In Fig. 2, we present experimental data from the channel illustrating the delay dependence of several key observables: (a) the angle between the momentum vectors of two O+ ions [], (b) the KE of S+ fragment [KE()], and (c) the total KE of two O+ fragments [KEsum()]. Simulation results are depicted in panels (d-f). The delay-dependent mean values of experimental data in (a-c) are shown in panel (g). Panel (h) features distributions of the ions at 250 fs (scatter plot) and 210 fs (contour line) when the bending wave packet is at the inner and outer turning points, respectively.
@@
There are two distinguished features in the experimental data shown in Fig. 2(a-c). First, all observables—, KE(), and KEsum()—show pronounced oscillations with a period of 83 fs (or 400 cm-1), clearly indicating that we mainly image the vibrational wave packet in the ground state of SO [70]. Second, there is a clear correlation between these observables: and KEsum() oscillate in-phase with each other and out-of-phase with KE(S+). This can be explained in terms of Coulomb forces between point charges (see Fig. S3 in the SM): near the equilibrium geometry, the net force on the S+ (O+) fragment due to the other two decreases (increases) with increasing bond angle. Moreover, also increases monotonically with the real-space bond angle.
We numerically model the experimental observables in Figs. 2(a-c) by solving the time-dependent Schrödinger equation for the coupled nuclear motion on the Born-Oppenheimer potential energy surfaces (BOS) of the SO2() and SO() states, including all vibrational degrees of freedom (symmetric stretch, bending, and antisymmetric stretch modes) [84]. We calculate the BOS ab initio by applying the multi-configurational self-consistent-field method as implemented in the quantum chemistry code GAMESS [85], based on seven frozen inner orbitals and 12 active orbitals expanded in the correlation consistent-polarized valence triple zeta (cc-pVTZ) basis set. We account for single ionization of SO2() to SO() by propagating the nuclear wave functions of the neutral and cationic states subject to dipole coupling by the pump pulse, with the assumption that the photoelectron has zero kinetic energy [84]. Within this ad hoc dipole-coupling model, the neutral ground state is depleted (and the ionic state populated) throughout the pulse, resulting in coherent vibrational wave packets in both states [75, 45]. To account for multiple ionization in the delayed probe pulse, we project the SO2 and SO nuclear wave functions onto the potential energy surface of the triply-charged molecule, approximated as purely Coulombic. We then propagate the ensuing nuclear wave packet representing the S+ and the two O+ ions to sufficiently large internuclear distances to reach numerical convergence of their momentum expectation values. We obtain the fragment KEs by FFT of the real-space nuclear wave function to momentum space [86].
Our simulations shown in Figs. 2(d-f) closely resemble the experimental data in Figs. 2(a-c), manifesting similar periodicity, phases, and correlations between the three observables. However, each pair of experiment-simulation plots shows a mismatch in absolute values of angle or KE. The primary reason for this mismatch is that vertical projection to and propagation on the Coulomb potential, which we used in our simulation, do not accurately describe the ionization and fragmentation process caused by the probe pulse. Furthermore, as will be discussed below, the wave packet propagation on the state, which extends to smaller bond angles and is not included in our simulation, also contributes to the experimental data (see also Fig. S3 in the SM). Nevertheless, our results provide a direct intuitive picture of the strong-field-induced vibrational wave packet motion in a triatomic molecule. Modulations of this three-body channel are dominated by the ionic ground-state vibrations, with a weak contribution from bending vibrations in the ground state of neutral SO2, as revealed by FFT of the delay-dependent data (Fig. 1 and SM, Fig. S2).
Besides revealing the states involved in the vibrational wave packet, coincidence momentum imaging also allows us to visualize the correlated motion of individual atoms. To achieve this, we plot the data for all ions from the channel in a 2D momentum image (Newton plot). Fig. 2(h) features a representative image at 250 fs with a contour for data at 210 fs (the inner and outer turning points of the bending wave packet). Because the and KE() manifest the same behavior as the OSO bond angle and the displacement of S from the center of mass (Fig. 2 and Fig. S3), movies comprising a timed sequence of these images (as provided in the SM) are a striking and intuitive representation of the bending vibration of the molecule. These molecular movies illustrate that during the first 400 fs, the wave packet is well localized and mimics the motion of a “ball-and-stick” model.
@@
While bending vibrations are responsible for the most pronounced features of the observables shown in Figs. 1 and 2, the experimental data contain more information about the molecular wave packet launched by the pump pulse. One hint of other dynamics is found in the total yield of the (O+, O+, S+) channel, which does not exhibit oscillations at the frequency of the bending vibration. This suggests that double ionization of the cation is insensitive to this mode. However, as depicted in Fig. 3(a), the yield initially decreases by approximately 15% within the first few hundred femtoseconds before flattening out. To explain this behavior, we also plot for this channel in Fig. 3(a). Here, is the angle between the laser polarization and , representing the direction of the O-O axis. The strong correlation between these two quantities (except near the overlap region) indicates that the observed yield variation is due to the initial alignment of the pump-induced rotational wave packet. The time scale is consistent with the rotational dynamics of the two-body breakup channels shown in Fig. S4 (SM). Although the O-O axis is the most polarizable axis of the molecule, strong-field ionization of neutral SO2 is more likely to occur along the symmetry axis (see [61] and SM, Fig. S5). Thus, SO ions are most likely created near the peak of the pump pulse with their symmetry axes aligned with the laser polarization. They then rotate due to the angular momentum accumulated in the pump pulse and reach peak alignment (of the O-O axis) around 200 fs. After the peak, the rotational wave packet dephases and exhibits long-term incoherent alignment.
This rovibrational wave packet motion in SO has other manifestations that can be elucidated by multi-coincidence momentum imaging. One example is the formation of O, which can be traced by analyzing the (S+,O) coincidence channel. The delay-dependent yield of this channel is shown in Fig. 3(b) along with the mean value of in the three-body channel. Clearly, the two oscillate out of phase, suggesting an intuitive picture of O formation by the probe pulse: O is more likely to be formed if the probe pulse finds the cation at smaller OSO angles, where the two oxygen atoms are closer to each other. The alignment and bending angle dependence of O production may help improve our understanding of a potential source of abiotic oxygen in SO2-rich planetary atmospheres [65].
The rovibrational wave packet also controls the excitation of the ground-state cations to the first excited state in the probe pulse. This single-photon transition is energetically allowed for and its dipole moment is along the O-O axis. Thus, when ground-state cations have rotated into alignment with the probe laser polarization and bending mode vibration has pushed the nuclear probability density to smaller-then-equilibrium bond angles (), a single photon in the leading edge of the probe can resonantly transfer population from to (Fig. 1). These two states have a symmetry-allowed conical intersection (CI) at with the branching space comprised of the bending and asymmetric stretching coordinates [60]. After excitation to the wave packet travels towards the CI where it is nonadiabatically coupled to the asymmetric stretch mode and returns to within 20 fs [60]. Since our probe pulse width is 28 fs, these cations can still be ionized at the peak of the probe and appear in the (O+, O+, S+) channel. On the other hand, if the excitation occurs in the trailing edge of the pump pulse, the wave packet should dephase rapidly as it traverses the CI and not exhibit 400 cm-1 oscillations.
In the experimental data, signatures of a wave packet in are seen in the KE sharing plot for the O+ ions. We separate events with and in Fig. 4(a,b), respectively, with the expectation that the population excited to can travel to smaller angles than the inner turning point of the bending wave packet in , but not to the outer turning point, within the probe pulse. The strong component with asymmetric KE sharing seen in Fig. 4(b) but not in Fig. 4(a) confirms the dipole coupling to and nonadiabatic coupling to the asymmetric stretch mode. Fig. 4(c) shows the delay-dependent yields of events with symmetric (blue) and asymmetric (green) KE sharing [all with small ]. The in-phase oscillation of the two curves is consistent with population transfer by the probe to near the inner turning point of the ground state bending vibration. To further test this hypothesis of resonant inter-state coupling, we use a single laser pulse at 405 nm to Coulomb-explode SO2 into (O+, O+, S+). At this shorter wavelength, efficient dipole coupling can occur only near the outer turning point of the wave packet and is unlikely to occur within our 30-fs pulse. The absence of asymmetric KE sharing peaks in the 405-nm case [Fig. 4(d)] confirms a resonant coupling near 790 nm, corroborating our conclusion that we see signatures of mode-coupling at the CI in our pump-probe experiment.
In conclusion, our work highlights the direct sensitivity of CEI to changes in the spatial density of the nuclear wave packet propagating on multiple electronic states, demonstrating the feasibility of tracking correlated atomic motion within a molecule. Our multidimensional coincident measurement allows us to observe and interpret the interplay between rotational, vibrational, and electronic motion, which determines the molecular wave packet evolution and experimental observables. Notably, correlated fragment momenta reveal clear signatures of a conical intersection that couples the bending and asymmetric stretching coordinates — which is difficult to access with other experimental techniques. This study paves the way for tracking coupled electronic and nuclear dynamics in medium-size polyatomic molecules, where high-order correlations revealed by CEI provide unique insights into the 3D correlated motion of multiple atomic constituents. Recent advancements have shown that CEI can image detailed 3D structures of gas-phase molecules with about ten atoms [25, 26, 87, 88], bringing this goal closer to reality.